Solve the equation for X
2/5 = x-2/20
To get this equation forst I must figure out what x is:
5-2=3
10-2=8
12-2=10
2/5 = x-2/20
2/5 = 10-2/20
Simplest form=
10/5=2
20/5=2
Leaving the eqation with
x=12
Is this right:

Answer by checkley71(8403)

2/5=X-2/20 THE EASIEST WAY TO SOLVE THIS EQUATION IS TO CONVERT ALL NUMBERS TO DECIMAL .
THUS 2/5=0.4 & 2/20=0.1
0.4=X-0.1
X=0.4+0.1
X=0.5 OR 1/2
PROOF
2/5=1/2-2/20
2/5=(10-2)/20
2/5=8/20 CROSS MULTIPLY
40=40
---------------------------------------------------
IF YOU MUST WORK IN FRATIONS THEN
2/5=X-2/20
X=2/5+2/20 FIND A COMMON DENOMINATOR (20)
X=(8+2)/20
X=10/20
X=1/2
###########################################
MERRY CHRISTMAS